OpenStudy (lena772):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

hartnn (hartnn):

can you find slope of y= -x+6 first ?

hartnn (hartnn):

compare it with y=mx+x m= slope

hartnn (hartnn):

**y= mx+c

OpenStudy (lena772):

-1

hartnn (hartnn):

correct. now product of slopes of perpendicular lines = -1

hartnn (hartnn):

so, m (-1) = -1 m= slope of required line = ... ?

OpenStudy (lena772):

So it's either A or C

OpenStudy (lena772):

-1

hartnn (hartnn):

A,C are parallel lines because they have same slope we need perpendicular line

OpenStudy (lena772):

Woops, my bad.

hartnn (hartnn):

no... m (-1) = -1 you got this, right ?

OpenStudy (lena772):

Oh so m needs to be 1

hartnn (hartnn):

yup!

hartnn (hartnn):

now use point slope form know what it is ?

OpenStudy (lena772):

y-y1=m(x-x1) right?

hartnn (hartnn):

correct

OpenStudy (lena772):

y--2=1(-1--4)

OpenStudy (lena772):

y+4=-1+4

hartnn (hartnn):

where did x go ?

OpenStudy (lena772):

Dangnabbit

OpenStudy (lena772):

y+4=x+2

hartnn (hartnn):

lol no :P

OpenStudy (lena772):

OMG. -_-

hartnn (hartnn):

y +2 = 1 (x+4)

OpenStudy (lena772):

Sigh, y+2=x+4

OpenStudy (lena772):

y=x+2

hartnn (hartnn):

\(\huge \color{red } \checkmark \)